OpenStudy (anonymous):

what is the slope of each of the lines. -4x+3y=3 5x+8y=0

OpenStudy (tkhunny):

Have you considered solving for y, putting them in Slope-Intercept form?

OpenStudy (anonymous):

i got this one wrong thats y i am asking for help to see other ways to solve it.

OpenStudy (anonymous):

@mathmale

OpenStudy (tkhunny):

Well, then solve for y and get it right. There is nothing stopping you but your algebra. Please show your work,

OpenStudy (nikato):

When trying to put into slope intercept form, you want to try to isolate y so it's y=..... It'll look like this: y=mx+b. where m=slope I'll show you the first and you can try the second on your own. -4x+3y=3 +4x. +4x ----------------- 3y=4x+3 -- -------- 3. 3 y=4/3x +3/3 y=4/3x +1 Using y=mx+b with m=slope You can see that 4/3 is the slope of this equation